OpenStudy (anonymous):

HELP WILL GIVE MEDALS!!! A bag of candy contains 3 blue, 5 red, 4 green, and 3 brown pieces.

OpenStudy (anonymous):

what are you, 5? this is easy

OpenStudy (anonymous):

a. What is the probability of picking a blue and then a green, with replacement? b. What is the probability of picking a blue and then a green, without replacement? c. The probability of choosing a blue, putting it back and picking another candy is 1/15. What was the second color picked? d. You have a new bag of 20 pieces of candy with the same colors of candy, just different amounts. You know there are 5 reds. You pick a red, don't put it back and then pick a green. The probability of this happening is 5/38. How many greens are there?

OpenStudy (anonymous):

actual question^

OpenStudy (asad786):

lol

OpenStudy (anonymous):

its algebra ll

OpenStudy (welshfella):

probability of picking a green = number of green pieces ---------------------- total number of pieces in the box

OpenStudy (welshfella):

for a blue its number of blue on top of fraction

OpenStudy (welshfella):

now for a blue and then a green you multiply the probabilities because both events are independent

OpenStudy (welshfella):

i this is the info for question a

OpenStudy (anonymous):

so and means multiply and or means add?

OpenStudy (welshfella):

yes

OpenStudy (anonymous):

ok

OpenStudy (welshfella):

P (green) = 3/15 right?

OpenStudy (anonymous):

its 4

OpenStudy (welshfella):

- did you spot my deliberate mistake??

OpenStudy (welshfella):

you did! lol

OpenStudy (anonymous):

yeah green is 4/15 not 3/15

OpenStudy (anonymous):

lol now so if its 3/15 , 4/14 i would multiply both denom and numerator

OpenStudy (welshfella):

yea - I made a mistake - it wasnt deliberate...

OpenStudy (anonymous):

but if it denominator is the same do i still multiply them

OpenStudy (welshfella):

4/15 not 4/14 ytes you multiply 3/15 * 4/15 or easier 1/5 * 4/15

OpenStudy (welshfella):

yes - its multiplication not addition

OpenStudy (anonymous):

well A says with replacement and B says without so A it would be 4/14 wouldnt it or is it B for 4/14

OpenStudy (welshfella):

Oh yes - i thought we were still talking about A

OpenStudy (anonymous):

wait so its B that turns into 4/14

OpenStudy (welshfella):

4/14 for B

OpenStudy (anonymous):

lol ok

OpenStudy (anonymous):

now for C the wording is confusing for me

OpenStudy (welshfella):

yes - first you pick out a blue put it back and then pick another candy but they dont tell you which one so you can call the second candy x and form an equation

OpenStudy (welshfella):

so (1/5) x = (1/15)

OpenStudy (anonymous):

ok so d would just be 3

OpenStudy (anonymous):

yeah i got that thanks

OpenStudy (welshfella):

so what would c be?

OpenStudy (anonymous):

i put that exact formula lol

OpenStudy (welshfella):

from the equation x = 1/15 / 1/5 = 1/15 * 5 = 1/3

OpenStudy (welshfella):

- so you need to find the candy whose probability of being picked is 1/3

OpenStudy (anonymous):

ok

OpenStudy (welshfella):

to make it easier write 1/3 as 5/15

OpenStudy (anonymous):

so would it be red

OpenStudy (welshfella):

right!!

OpenStudy (anonymous):

wait i think im wrong

OpenStudy (anonymous):

oh damn lol

OpenStudy (anonymous):

it looked wrong for a sec

OpenStudy (welshfella):

no - there are 5 reds Red is correct

OpenStudy (anonymous):

but yeah so d, i would its 3

OpenStudy (anonymous):

oh damn never mind.

OpenStudy (anonymous):

oh damn idk. them words

OpenStudy (welshfella):

hhmm P (Red) = 5/20 = 1/4 so (1/4) * x = 5/38

OpenStudy (anonymous):

multiply 4 by 1/4 ?

OpenStudy (welshfella):

its easier to write this as follows then cross multiply